# soft open fences

### v2

@@ Appendix B More Related Work
Appendix C Experimental Settings
C.1 Checkpoints
Usage Models Links
Llama-3.2-1B https://huggingface.co/meta-llama/Llama-3.2-1B
OpenAI GPT-4o https://platform.openai.com/docs/models/gpt-4o
o1-mini https://platform.openai.com/docs/models/o1
Table 3: Checkpoints used in experiments and their links.
C.2 Hyperparameters
Hyperparameter Value
temperature 1.0
top-k 1.0
top-p 1.0
repetition_penalty 1.0
max_new_tokens 200
max_seq_len 32768
[…]
Table 4: LLM Hyperparameters and EOS tokens used in experiments.
@@ Appendix D Algorithm Details of SC-MCTS
Appendix E Reward Functions Correlation
Figure 7: Reward Functions Correlation Heatmap.
@@ F.2 Prompts Settings of Easy Blocksworld
Input Instructions:
[STATEMENT]
As initial conditions I have that, the red block is clear, the hand is empty, the blue block is on top of the orange block, the red block is on the table, the orange block is on the table and the yellow block is on the table.
My goal is to have that the orange block is on top of the blue block. My plan is as follows:
[End Of STATEMENT]
[PLAN]
unstack the blue block from on top of the orange block
put down the blue block
pick up the orange block
stack the orange block on top of the blue block
[PLAN END]
[STATEMENT]
[End Of STATEMENT]
Output format:
[PLAN]
[LLM Completion]
[PLAN_END]
@@ F.3 Prompts Settings of Hard Blocksworld
Appendix G OpenAI API Data
@@ Appendix H GPU Usage
Appendix I Future Work

### v3

@@ Appendix B More Related Work
Appendix C Reward Functions Correlation
Figure 7: Reward Functions Correlation Heatmap.
@@ Appendix D Algorithm Details of SC-MCTS
Appendix E Experimental Settings
E.1 Checkpoints
Usage Models Links
Llama-3.2-1B https://huggingface.co/meta-llama/Llama-3.2-1B
OpenAI GPT-4o https://platform.openai.com/docs/models/gpt-4o
o1-mini https://platform.openai.com/docs/models/o1
Table 3: Checkpoints used in experiments and their links.
E.2 Hyperparameters
Hyperparameter Value
temperature 1.0
top-k 1.0
top-p 1.0
repetition_penalty 1.0
max_new_tokens 200
max_seq_len 32768
[…]
Table 4: LLM Hyperparameters and EOS tokens used in experiments.
@@ F.2 Prompts Settings of Easy Blocksworld
Input Instructions: I am playing with a set of blocks where I need to arrange the blocks into stacks. Here are the actions I can do: 1. Pick up a block 2. Unstack a block from on top of another block 3. Put down a block 4. Stack a block on top of another block I have the following restrictions on my actions: 1. I can only pick up or unstack one block at a time. 2. I can only pick up or unstack a block if my hand is empty. 3. I can only pick up a block if the block is on the table and the block is clear. A block is clear if the block has no other blocks on top of it and if the block is not picked up. 4. I can only unstack a block from on top of another block if the block I am unstacking was really on top of the other block. 5. I can only unstack a block from on top of another block if the block I am unstacking is clear. Once I pick up or unstack a block, I am holding the block. 1. I can only put down a block that I am holding. 2. I can only stack a block on top of another block if I am holding the block being stacked. 3. I can only stack a block on top of another block if the block onto which I am stacking the block is clear. Once I put down or stack a block, my hand becomes empty. [STATEMENT] As initial conditions I have that, the red block is clear, the hand is empty, the blue block is on top of the orange block, the red block is on the table, the orange block is on the table and the yellow block is on the table. My goal is to have that the orange block is on top of the blue block. My plan is as follows: [End Of STATEMENT] [PLAN] unstack the blue block from on top of the orange block put down the blue block pick up the orange block stack the orange block on top of the blue block [PLAN END] [STATEMENT] As initial conditions I have that, the red block is clear, the yellow block is clear, the hand is empty, the red block is on top of the blue block, the yellow block is on top of the orange block, the blue block is on the table and the orange block is on the table. My goal is to have that the orange block is on top of the red block. My plan is as follows: [End Of STATEMENT] Output format: [PLAN] [LLM Completion] [PLAN_END]
@@ F.3 Prompts Settings of Hard Blocksworld
Appendix G Example Trees of Different $c$ of UCT
Figure 8: Monte Carlo Tree with origin parameter $c$ of UCT
Figure 9: Monte Carlo Tree with our optimized parameter $c$ of UCT
From Figure 8 and 9 we can observed that with our optimized parameter $c$ of UCT, MCTS algorithm in node selection decisions tends to prioritize exploring new nodes rather than repeatedly following old paths, which may often lead to dead ends.
Appendix H OpenAI API Data
@@ Appendix I GPU Usage
Appendix J Future Work